OpenStudy (anonymous):

You flip three coins. What is the probability that you get at least two tails, given that you get at least one tail?

OpenStudy (amistre64):

um .... what?

OpenStudy (amistre64):

3 coins gives the sample space of: hhh hht tth hth tht thh htt ttt or 8 total

OpenStudy (anonymous):

sample space is thh hth hht tth tht htt ttt since this are the ones with at least one tail. of those 4 have at least 2 tails so answer is 4/7

OpenStudy (amistre64):

P(<=2tails) = 4/8 P(<=1 tail) = 7/8 ... maybe

OpenStudy (amistre64):

P(AnB) = 4/8; so maybe if i read it right: 4/8 --- = 4/7 ? 7/8

OpenStudy (anonymous):

hello my guru. it says "given that you have at least one tail" which is a conditional probability. you already know one of them is tails, so hhh is out. that is conditional sample space has 7 elements not 8. of the 7 4 have at least two tails, so get 4/7

OpenStudy (anonymous):

if you want a formula you can use \[P(A|B)=\frac{P(A\cap B)}{P(B)}=\frac{\frac{4}{8}}{\frac{7}{8}}\]

OpenStudy (amistre64):

thats what I used :)

OpenStudy (anonymous):

thank u cuz i have no clue what im doin here

OpenStudy (amistre64):

just wasnt sure if i was reading the ? right :)

OpenStudy (anonymous):

ah i see. you are right. site is slow slow slow. or else it is my computer